# Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Fine-tuning

### v4

@@ 4.2 AEPO
REINFORCE as regularization. In RLVR, the reward space is binary. This property allows REINFORCE to filter out negative samples in an unbiased manner, ensuring that the gradient is formed from positive samples that align with the desired distribution. Consequently, AEPO’s regularization term produces a one-sided optimization signal that guides the policy toward higher-quality behavior distributions.
Table 2: Main results on mathematical reasoning benchmarks. We compare Qwen2.5-math-7B (base model), GRPO, entropy-based baselines, and AEPO under different entropy thresholds $\mathcal{H}$. AEPO consistently outperforms all baselines, achieving stable entropy control and delivering substantial improvements across diverse benchmarks, with the best overall performance at $\mathcal{H}=0.75$.
Benchmarks AIME24 AMC College Math GSM8K MATH Minerva Math Olympiad Average
Qwen2.5-math-7B 13.3 40.0 41.7 65.4 65.5 11.0 26.7 37.66
GRPO 36.7 75.0 48.2 88.9 80.5 34.6 41.8 57.96
Entropy-Reg 36.7 75.0 47.6 87.0 80.4 35.7 40.4 57.39
Entropy-Adv 36.7 75.0 47.8 87.8 80.4 37.5 42.1 58.18
AEPO $\mathcal{H}=0.25$ 40.0 77.5 48.5 89.4 80.6 33.5 42.2 58.81
AEPO $\mathcal{H}=0.50$ 43.3 82.5 48.8 89.5 81.6 38.2 43.0 60.99
AEPO $\mathcal{H}=0.75$ 50.0 (+36.7) 80.0 (+40.0) 48.2 (+6.50) 89.4 (+24.0) 82.0 (+16.5) 37.5 (+26.5) 42.4 (+15.7) 61.36 (+23.70)
AEPO $\mathcal{H}=1.00$ 33.3 75.0 48.4 88.7 80.8 37.9 42.1 58.03
@@ 5 Experiments
To validate the effectiveness of our methods, we present the experimental setup and results in the following sections. Our work is based on the EasyR1 and VeRL frameworks (Yaowei et al., 2025; Sheng et al., 2025), and we compare with the RL baselines GRPO (Shao et al., 2024) and its entropy-regularization variants (Hou et al., 2025; Cheng et al., 2025).
@@ 5.1 Experimental setup
Model and Dataset: We conduct experiments to evaluate the effectiveness of AEPO in RFT for mathematical reasoning tasks. The base model is Qwen2.5-Math-7B (Yang et al., 2024), an LLM specialized for mathematical problem solving. For training, we use the DAPO-17K dataset (Yu et al., 2025), which contains diverse problem instances curated for RFT.
Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Tang et al., 2024), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023). These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
Implementation: We initialize all policy models and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout group size GG=5, temperature 1.0, learning rate $10^{-6}$, $\epsilon=0.2$, and a binary reward. For AEPO, $T_{\text{high}}=1.2$, $T_{\text{low}}=0.8$, we replace 60 temperature-adjusted positive samples in each batch. For entropy-regularized baselines, we set $\lambda=0.03$ for Entropy-Reg, and $\beta=0.4$, $\kappa=2$ for Entropy-Adv (Cheng et al., 2025). To exclude potential confounding factors, we did not apply KL divergence in any of our experiments.
@@ 5.2 Main results
Table 2 summarizes the performance of AEPO compared with GRPO and entropy-based baselines across seven mathematical reasoning benchmarks. We observe that AEPO consistently outperforms all competing methods, achieving the best results on every benchmark. In particular, AEPO yields an average score of 61.36, representing a relative improvement of +3.40 points over GRPO. Notably, the gains are especially pronounced on high-difficulty datasets such as AIME24 and AMC, where exploration plays a critical role in finding correct reasoning paths.
Compared to entropy-regularized variants, AEPO also exhibits clear advantages. Both Entropy-Reg and Entropy-Adv deliver only marginal improvements over GRPO. As shown in Fig. 4, the entropy regularization term in Entropy-Reg introduces significant bias: during mid-stage training, entropy begins to surge uncontrollably, indicating that entropy gradually replaces accuracy as the dominant optimization signal, with accumulated bias distorting the learning process. Entropy-Adv, on the other hand, fails to fundamentally reverse the trend of entropy collapse, leaving the model with insufficient exploration. In contrast, AEPO achieves consistent improvements across all benchmarks, underscoring the benefits of principled entropy control over heuristic entropy bonuses.
Figure 4: Entropy trajectories of AEPO compared with GRPO, Entropy-Reg, and Entropy-Adv. While GRPO and Entropy-Adv suffer from monotonic entropy collapse and fail to maintain exploration, Entropy-Reg leads to unstable fluctuations. In contrast, AEPO stabilizes entropy around the target level, demonstrating controllable and robust entropy regulation throughout training.
5.3 Entropy and exploration
Table 3: Comparison of different ablation losses: entropy control ability, and benchmark performance.
Loss and performance Entropy control
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$
AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
33.3 75.0 47.7 88.9 80.0 33.5 40.1 56.93 (-4.43)
entropy collapse
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$
AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
36.7 75.0 48.3 88.1 80.8 36.0 41.8 58.14 (-3.22)
entropy collapse
To further examine the relationship between entropy and exploration, we evaluate AEPO under different entropy thresholds $\mathcal{H}$. As shown in Table 2, AEPO consistently surpasses GRPO across all benchmarks regardless of the entropy target, with strong improvements on challenging datasets such as AIME24 and MATH. At different thresholds, AEPO maintains clear advantages, indicating that principled entropy regulation yields benefits across a wide range of entropy levels.
Figure 1 further highlights this effect. While GRPO exhibits monotonic entropy collapse, AEPO stably maintains entropy around arbitrary preset levels simply by adjusting the hyperparameter $\mathcal{H}$, since AEPO is not sensitive to the parameter $\alpha$, as long as it is chosen within a reasonable range. This controllability allows us to probe the direct impact of entropy on exploration and downstream performance. Notably, we observe a non-monotonic trend: performance first improves as entropy rises from low to moderate levels, but begins to decline when entropy becomes excessively high. This provides explicit evidence that moderate entropy fosters effective exploration, while overly high entropy disperses optimization and harms accuracy.
Together, these results confirm that AEPO not only eliminates entropy collapse but also clarifies the link between entropy, exploration, and performance, offering a controllable pathway to balance exploration with convergence.
@@ 6 Ablation Study
To assess the contribution of each component in AEPO, we design two ablation studies to verify the necessity of distribution as regularization and REINFORCE as regularization for achieving effective entropy control.
6.1 Distribution as regularization
One critical component of AEPO is the use of temperature-adjusted distributions for entropy control. In our design, the REINFORCE regularization term samples from a modified distribution, where the temperature $T$ is adaptively adjusted based on the previous step’s entropy. This adjustment ensures that positive samples carry either higher or lower entropy as required, thereby stabilizing the overall entropy around the target threshold $\mathcal{H}$.
To validate the necessity of this design, we replace the temperature-adjusted distribution with the original distribution, as shown in Table 3. The results demonstrate that when REINFORCE samples are drawn directly from the original policy distribution (i.e., without temperature adjustment), entropy control collapses: the policy entropy monotonically decreases during training, similar to GRPO. More importantly, the average benchmark score drops to 57.39, which is even worse than standard GRPO and far below AEPO (61.36). This shows that the variance of REINFORCE gradients, when unregularized by distribution adjustment, further degrades optimization stability and downstream reasoning accuracy.
These findings confirm that distribution adjustment is indispensable for AEPO: it directly enables controllable entropy regulation, reduces variance, and consistently improves reasoning performance across benchmarks. Moreover, it provides strong evidence for the role of policy gradient as regularization: the REINFORCE term in AEPO influences the exploration ability of the GRPO component through entropy control, thereby shaping the overall optimization dynamics during training.
@@ 6.2 REINFORCE as regularization
To examine the role of REINFORCE, we conduct an ablation where the regularization term is still sampled from temperature-adjusted distributions, but the advantage function $\hat{A}_{t}$ is used directly without filtering negative samples. As shown in Table 3, this variant fails to prevent entropy collapse, leading to degraded performance across benchmarks. The average score drops to 58.14. This confirms that filtering negative samples via REINFORCE is indispensable.
@@ 7 Conclusion and discussion
In this paper, we propose Arbitrary Entropy Policy Optimization (AEPO), a principled RFT framework that addresses one of the most persistent challenges in RFT—precise and stable entropy control. Unlike traditional entropy regularization methods that trade off exploration against stability, AEPO achieves controllable entropy regulation through a unified design that integrates policy gradient, distribution, and REINFORCE as regularization components. This formulation eliminates the entropy collapse phenomenon in GRPO and maintains policy entropy within an arbitrarily specified range, enabling balanced and consistent exploration throughout training.
Extensive experiments across seven mathematical reasoning benchmarks demonstrate that AEPO consistently outperforms entropy-based baselines, exhibiting greater stability, generalization, and robustness to hyperparameters. More importantly, AEPO reveals a non-monotonic relationship between entropy and reasoning performance, showing that moderate entropy fosters exploration while excessive entropy impairs optimization—offering the first quantitative evidence linking entropy dynamics to reasoning capability in large language models.
Beyond solving entropy collapse, AEPO establishes a generalizable paradigm for learning under target distributions, suggesting its applicability to broader domains such as multi-modal alignment and long-horizon reasoning. In summary, AEPO transforms entropy from a passive statistical indicator into an active, tunable variable that fundamentally governs exploration, stability, and performance in RFT.

### v6

@@ 4.2 AEPO
REINFORCE as regularization. In RLVR, the reward space is binary. This property allows REINFORCE to filter out negative samples in an unbiased manner, ensuring that the gradient is formed from positive samples that align with the desired distribution. Consequently, AEPO’s regularization term produces an unidirectional optimization signal that guides the policy toward higher-quality behavior distributions.
Figure 4: Entropy trajectories of AEPO compared with GRPO, Entropy-Reg, and Entropy-Adv. AEPO stabilizes entropy around a moderate level, demonstrating controllable and robust entropy regulation throughout training.
Table 1: AEPO demonstrates consistently superior performance across all mathematical reasoning benchmarks, surpassing GRPO, and all entropy-regularized baselines.
Benchmarks AIME24×32 AIME25×32 AMC×32 GSM8K MATH Minerva Olympiad Average
Qwen2.5-7B 7.91 5.31 36.2 88.5 64.4 22.0 29.3 36.24
+GRPO 17.1 7.60 65.8 92.3 75.6 36.8 38.8 47.70
+Entropy-Reg 13.6 8.85 67.4 92.3 76.8 35.5 39.1 47.65
+Entropy-Adv 14.8 8.23 67.3 91.9 76.6 38.2 37.5 47.79
+AEPO    $\Delta$ vs. GRPO 17.5 (+0.4) 11.4 (+3.8) 69.3 (+3.5) 92.9 (+0.6) 78.0 (+2.4) 37.8 (+1.0) 40.2 (+1.4) 49.57 (+1.87)
Qwen2.5-math-7B 15.5 7.81 42.1 65.4 59.4 11.0 26.7 32.56
+GRPO 32.1 11.0 72.4 88.7 80.6 34.6 41.8 51.60
+Entropy-Reg 31.4 10.1 74.3 87.0 78.8 35.7 40.4 51.10
+Entropy-Adv 32.1 11.4 72.1 87.8 78.8 37.5 42.1 51.76
+AEPO    $\Delta$ vs. GRPO 36.4 (+4.3) 12.6 (+1.6) 74.8 (+2.4) 89.5 (+0.8) 82.6 (+2.0) 38.2 (+3.6) 43.0 (+1.2) 53.87 (+2.27)
Qwen3-4B 36.4 22.7 71.9 93.9 84.8 42.3 47.2 57.03
+GRPO 52.9 41.5 86.1 95.2 92.0 46.7 60.0 67.77
+AEPO    $\Delta$ vs. GRPO 54.5 (+1.6) 43.7 (+2.2) 89.7 (+3.6) 95.0 (-0.2) 92.8 (+0.8) 47.8 (+1.1) 60.9 (+0.9) 69.20 (+1.43)
Table 2: Comparison of pass@$k$ performance between the base model, GRPO, and AEPO across four mathematical reasoning benchmarks (steps $\geq$ 100). The accompanying figures show the evolution of pass@1024 of AEPO and GRPO on AIME24 and AIME25 during training.
Benchmarks AIME24 AIME25 Minerva Olympiad
Pass@512 Pass@1024 Pass@512 Pass@1024 Pass@128 Pass@128
Qwen2.5-7B 71.7 80.0 61.7 73.3 69.8 77.8
+GRPO 63.3 73.3 60.0 63.3 65.8 72.9
+AEPO 70.0 83.3 66.7 73.3 67.0 74.1
Table 3: Ablation study on AEPO under different entropy targets, including benchmark performance across seven mathematical reasoning datasets. The upper table compares AEPO at various entropy levels, while the lower ablations illustrate how different loss formulations affect both performance and entropy stability.
Benchmarks AIME24×32 AIME25×32 AMC×32 GSM8K MATH500 Minerva Olympiad Average
Qwen2.5-math-7B 15.5 7.81 42.1 65.4 59.4 11.0 26.7 37.66
+AEPO $\mathcal{H}=0.25$ 37.9 11.3 74.4 89.4 79.8 36.0 42.2 53.00
+AEPO $\mathcal{H}=0.50$ 36.4 12.6 74.8 89.5 82.6 38.2 43.0 53.87
+AEPO $\mathcal{H}=0.75$ 33.6 15.1 77.0 89.4 79.2 37.5 42.4 53.45
+AEPO $\mathcal{H}=1.00$ 33.2 15.6 74.8 88.7 79.6 37.9 42.1 53.13
Ablation loss and performance Entropy control
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$    (2)
AIME24 AIME25 AMC GSM8K MATH500 Minerva Olympiad Avg
31.4 11.8 74.5 88.9 79.0 33.5 40.1 51.31 (-2.56)
Entropy collapse
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$    (3)
AIME24 AIME25 AMC GSM8K MATH500 Minerva Olympiad Avg
32.7 11.0 73.2 88.1 79.0 35.7 41.2 51.56 (-2.31)
Entropy collapse
@@ 5 Experiments
To validate the effectiveness of our methods, we present the experimental setup and results in the following sections. Our work is based on the EasyR1 and VeRL frameworks (zheng2025easyr1; sheng2025hybridflow), and we compare with the RL baselines GRPO (shao2024deepseekmath) and its entropy-regularization variants (hou2025advancing; cheng2025reasoning).
@@ 5.1 Experimental setup
Model and Dataset: We conduct experiments to evaluate the effectiveness of AEPO in RL for mathematical reasoning tasks. The base models include Qwen2.5-7B, Qwen2.5-Math-7B and Qwen3-4B. (yang2024qwen2; yang2025qwen3). For training, we use the DAPO-17K dataset (yu2025dapo), which contains diverse problem instances curated for RL.
Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24, AIME25 (hf_aime2024), AMC (lightman2023lets), GSM8K (cobbe2021gsm8k), MATH (lightman2023lets), Minerva Math (lewkowycz2022solving), and Olympiad (lightman2023lets). These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
@@ 5.2 Main results
Breaking the entropy bottleneck. Fig. 1 and 4 show that AEPO fundamentally resolves the entropy collapse issue that hampers GRPO and other RL baselines: AEPO maintains entropy around arbitrary target levels and avoids both collapse and uncontrolled expansion. As summarized in Table 1, on almost all mathematical reasoning benchmarks, AEPO surpasses GRPO and entropy-based baselines, demonstrating a much stronger capacity for effective exploration. In contrast, entropy-regularized baselines either fail to sustain entropy or introduce significant optimization bias that disrupts learning.
Breaking the exploration bottleneck. AEPO delivers consistent improvements over RL baselines on both pass@1 and pass@$k$ across every benchmark. More importantly, as shown in the Table. 2, AEPO shows the ability to explore broader than the base model on pass@1024, demonstrating that RL can meaningfully expand the reasoning frontier rather than merely sharpening the base model’s existing knowledge. During training, the pass@1024 performance of GRPO decreases monotonically, whereas AEPO shows the ability to expand the reasoning frontier, exhibiting upward trends rather than a purely decreasing curve. These results confirm that exploration can genuinely unlock new reasoning behaviors that are inaccessible to standard RL methods when guided by principled entropy modulation.
Entropy-performance relation. By varying the entropy target, AEPO enables a controlled study of how entropy affects exploration and downstream accuracy. Across all entropy levels, AEPO consistently outperforms GRPO, demonstrating that stable entropy regulation is inherently beneficial. More interestingly, as shown in Table 3, AEPO reveals a clear non-monotonic relationship: moderate entropy improves exploration and enhances performance, whereas overly high entropy disperses optimization and reduces accuracy. Different benchmarks also exhibit varying degrees of benefit from exploration; some continue to gain as entropy increases, while others experience performance degradation. Overall, the results display a rise-then-fall trend, indicating the presence of an optimal entropy regime for effective reasoning.
@@ 6 Ablation Study
To assess the contribution of each component in AEPO, we design two ablation studies to verify the necessity of temperature as regularization and REINFORCE as regularization for achieving effective entropy control.
6.1 Temperature as regularization
One critical component of AEPO is the use of temperature-adjusted samples for entropy control. In our design, the REINFORCE regularization term samples from a modified distribution, where the temperature $T$ is adaptively adjusted based on the previous step’s entropy. This adjustment ensures that positive samples carry either higher or lower entropy as required, thereby stabilizing the overall entropy around the target threshold $\mathcal{H}$.
To validate the necessity of this design, we replace the temperature-adjusted distribution with the original distribution, as shown in Eq. (3) of Table 3. The results demonstrate that when REINFORCE samples are drawn directly from the original policy distribution (i.e., without temperature adjustment), entropy control collapses: the policy entropy monotonically decreases during training, similar to GRPO. More importantly, the average benchmark score drops to 51.31, which is even worse than standard GRPO and far below AEPO (53.87). This shows that the variance of REINFORCE gradients, when unregularized by distribution adjustment, further degrades optimization performance.
These findings confirm that temperature adjustment is indispensable for AEPO: it directly enables controllable entropy regulation, and consistently improves reasoning performance across benchmarks. Moreover, it provides strong evidence for the relation between entropy and exploration: the REINFORCE term in AEPO influences the exploration ability of the GRPO component through entropy control, thereby shaping the overall optimization dynamics during training.
@@ 6.2 REINFORCE as regularization
To examine the role of REINFORCE, we conduct an ablation where the regularization term is still sampled from temperature-adjusted distributions, but the advantage function $\hat{A}_{t}$ is used directly without filtering negative samples. As shown in Eq. (3) of Table 3, this variant fails to prevent entropy collapse, leading to degraded performance across benchmarks. The average score drops to 51.56. This confirms that filtering negative samples via REINFORCE is indispensable.
@@ 7 Conclusion and discussion
In this paper, we propose AEPO, a principled RL framework that addresses one of the most persistent challenges in RL—precise and stable entropy control. Unlike traditional entropy regularization methods that trade off exploration against stability, AEPO achieves controllable entropy regulation through a unified design that integrates policy gradient, distribution, and REINFORCE as regularization components. This formulation eliminates the entropy collapse phenomenon in GRPO and maintains policy entropy within an arbitrarily specified range, enabling balanced and consistent exploration throughout training.
Extensive experiments across seven mathematical reasoning benchmarks demonstrate that AEPO consistently outperforms entropy-based baselines on both pass@1 and pass@k, exhibiting greater stability, generalization, and robustness to hyperparameters. Moreover, AEPO shows that RL-based exploration can indeed move beyond the limitations of the base model, confirming that reinforcement learning is capable of expanding the reasoning frontier rather than merely refining pretrained knowledge. More importantly, AEPO reveals a non-monotonic relationship between entropy and reasoning performance, showing that moderate entropy fosters exploration while excessive entropy impairs optimization—offering the first quantitative evidence linking entropy dynamics to reasoning capability in large language models.
Beyond entropy control, AEPO provides a broadly generalizable framework for learning under target distributions. The temperature-based mechanism used in AEPO should be viewed only as a special case: temperature creates the target distribution. When the task objective changes, the same principle naturally extends beyond entropy control. For example, if the goal is to mitigate overthinking by encouraging shorter responses, and if one can construct a transformed distribution $\pi^{*}$ from the current policy $\pi$ that produces more concise outputs while preserving correctness, then AEPO can use samples from $\pi^{*}$ to form a REINFORCE regularization. In doing so, the current policy is gradually guided toward acquiring the desired properties of $\pi^{*}$. More generally, whenever a target distribution with specific behavioral characteristics can be derived from the current policy, AEPO draws samples from this target and applies a REINFORCE-based regularization term, enabling the current policy to absorb the desired behavioral characteristics from the target distribution.
AEPO is not merely an entropy-control technique but a general paradigm for steering policies toward arbitrarily defined target behaviors through principled policy-gradient regularization. The same mechanism that adjusts the entropy can therefore be extended to new domains, illustrating the strong generalizability of AEPO.